# Geometric Height on flag varieties in positive characteristic

### v1

@@ 3.1. A height lower bound in Schubert cells
In subsection 3.1 we follow [3] since there is no difference between characteristic zero and characteristic $p$. For $w\in W_{Q}\backslash W/W_{P}$, write $\mathcal{C}_{w}=F_{Q}\times_{Q}QwP/P,\mathcal{X}_{w}=F_{Q}\times_{Q}\overline{QwP}/P,C_{w}=\mathcal{C}_{w,K}$ and $X_{w}=\mathcal{X}_{w,K}$.
Let $F_{Q}$ be the canonical reduction of $F$ to $Q$, and ${\mathrm{deg}}(F_{Q})\in X(T)^{\vee}_{\mathbb{Q}}$ be the degree cocharacter. Let $W,W_{P}$ and $W_{Q}$ be the Weyl group of $G,L_{P}$ and $L_{Q}$. For $w\in W_{Q}\backslash W/W_{P}$, write $\mathcal{C}_{w}=F_{Q}\times_{Q}QwP/P$, $\mathcal{X}_{w}=F_{Q}\times_{Q}\overline{QwP}/P$, $C_{w}=\mathcal{C}_{w,K}$ and $X_{w}=\mathcal{X}_{w,K}$.
Note that for any $w^{\prime}\in W_{Q}$ and $\lambda\in X(T)$, we have $w^{\prime}\lambda-\lambda\in\mathbb{Z}\left[\Delta_{Q}\right]$ and consequently $\left\langle\operatorname{deg}\left(F_{Q}\right),w^{\prime}\lambda\right\rangle=\left\langle\operatorname{deg}\left(F_{Q}\right),\lambda\right\rangle$. Note also that for any $\lambda\in X(P)$ and $w\in W_{P},w\lambda=\lambda$. So the number $\left\langle\operatorname{deg}\left(F_{Q}\right),w\lambda\right\rangle$ is well-defined for any $\lambda\in X(P)$ and $w\in W_{Q}\backslash W/W_{P}$.
@@
The proof in [3] does not require ${\rm char}(k)=0$, therefore works in our case. ∎
@@ 3.2. Height filtration and successive minima
Let $V$ be any $Q$-representation of highest weight $\lambda\in X^{*}(T)$ and let $V=\bigoplus_{\nu}V[\nu]$ be its weight decomposition. Furthermore, let $F_{Q}$ be the strongly canonical reduction of $F$. We define a filtration $V_{\bullet}$ on the vector space $V$: For any rational number $q\in\mathbb{Q}$, we define the subspace $V_{q}$ as the sum of weight spaces
@@
Then, by twisting the $Q$-subrepresentations $V_{q}$ above by $F_{P}$, we obtain a filtration $V_{\bullet,F_{Q}}$ of the vector bundle $V_{F_{G}}=F_{G}\times^{G}V$ by subbundles
$0\neq V\left[\lambda+\mathbb{Z}\Delta_{M}\right]_{F_{M}}=(\bigoplus_{\nu\in\lambda+\mathbb{Z}\Delta_{M}}V[\nu])_{F_{M}}=V_{q_{1},F_{Q}}\subsetneq\cdots\subsetneq V_{q,F_{Q}}\subsetneq\cdots\subsetneq V_{q_{0},F_{Q}}=V_{F_{G}}.$
Improving a result of Schieder [^1] The case of characteristic 00 and a specific case of positive characteristic were proven in [8], we proved that:
@@
We will need the following well-known theorem.
Theorem 3.6 ([6], Theorem 3.23).
@@
It will suffice to check two assertions: all qraded pieces are semistable vector bundles and their slopes are decreasing.
As shown in [8] Proposition 5.1, we have
$\mathrm{gr}_{q}(V_{\bullet,F_{P}})=(\mathrm{gr}_{q}V_{\bullet})_{F_{M}}.$
And its slope is computed as
$\mu((\mathrm{gr}_{q}V_{\bullet})_{F_{M}})=q.$
@@
To prove the semistability, consider the semisimplification $W$ of $(\mathrm{gr}_{q}V_{\bullet})$ as a representation of $M$. Now apply [8] Proposition 3.2(a), we see that for each simple direct summand $W_{i}$ of $W$ the associated bundle $W_{i,F_{M}}$ has slope $q$. Thus $(\mathrm{gr}_{q}V_{\bullet})_{F_{M}}$ admits a filtration by subbundles such that each graded piece is a semistable vector bundle of slope $q$, where the semistability is obtained by applying Theorem 3.6 to $M\to{\rm GL}(M_{i})$. Then $(\mathrm{gr}_{q}V_{\bullet})_{F_{M}}$ is also semistable of slope $q$ by the following. ∎
@@
If $E$ is a vector bundle on $C$ that admits a filtration of subbundles
@@
and all graded piece are semistable of slope equals to $\lambda$, so is $E$.
@@
The first assertion is obtained by taking $V$ to be $H^{0}\left(\overline{QwP}/P,M_{\lambda}\right)$ in Proposition 3.5. For the second assertion, it will suffice to show the highest weights in $H^{0}\left(\overline{QwP}/P,M_{\lambda}\right)$ belong to $w\lambda+\mathbb{Z}[\Delta_{Q}]$. The argument in [3] is also valid without any assumption on the characteristic of base field. Note that one can also apply [4, § 14 Proposition 12] in positive characteristic, since Schubert variety in positive characteristic is normal by [7]. ∎
@@
The arguments follows from [3] once the lower bound and essential minimum is obtained. See [3, Proof of Theorem 2.1]
@@
[…]

### v2

@@ 3.1. A height lower bound in Schubert cells
In subsection 3.1, we follow [4] since there is no difference between characteristic zero and characteristic $p$. For $w\in W_{Q}\backslash W/W_{P}$, write $\mathcal{C}_{w}=F_{Q}\times_{Q}QwP/P,\mathcal{X}_{w}=F_{Q}\times_{Q}\overline{QwP}/P,C_{w}=\mathcal{C}_{w,K}$ and $X_{w}=\mathcal{X}_{w,K}$.
Let $F_{Q}$ be the canonical reduction of $F$ with canonical parabolic $Q$, and ${\mathrm{deg}}(F_{Q})\in X(T)^{\vee}_{\mathbb{Q}}$ be the degree cocharacter. Let $W,W_{P}$ and $W_{Q}$ be the Weyl groups of $G,L_{P}$ and $L_{Q}$. For $w\in W_{Q}\backslash W/W_{P}$, write $\mathcal{C}_{w}=F_{Q}\times_{Q}QwP/P$, $\mathcal{X}_{w}=F_{Q}\times_{Q}\overline{QwP}/P$, $C_{w}=\mathcal{C}_{w,K}$ and $X_{w}=\mathcal{X}_{w,K}$.
Note that for any $w^{\prime}\in W_{Q}$ and $\lambda\in X(T)$, we have $w^{\prime}\lambda-\lambda\in\mathbb{Z}\left[\Delta_{Q}\right]$ and consequently $\left\langle\operatorname{deg}\left(F_{Q}\right),w^{\prime}\lambda\right\rangle=\left\langle\operatorname{deg}\left(F_{Q}\right),\lambda\right\rangle$. Note also that for any $\lambda\in X(P)$ and $w\in W_{P},w\lambda=\lambda$. So, the number $\left\langle\operatorname{deg}\left(F_{Q}\right),w\lambda\right\rangle$ is well-defined for any $\lambda\in X(P)$ and $w\in W_{Q}\backslash W/W_{P}$.
@@
We follow the argument in [4], since many techniques do not require the characteristic.
For $x\in X(K)$, let $\sigma_{x}:C\to F/P$ be the section induced by $x:\operatorname{Spec}(K)\to X$ through the valuative criterion. Let $F_{P,x}=\sigma_{x}^{*}F$ be the corresponding reduction to $P$.
Definition 3.4.
A reduction $F_{P}$ of $F$ to $P$ is called in relative position $w\in W_{Q}\setminus W/W_{P}$ with respect to $F_{Q}$ if the image of the natural map
$F_{Q}\times_{C}F_{P}\to G_{C},\quad(a,b)\mapsto a^{-1}b$
lies in $QwP_{C}$.
This definition coincides with the definition in [10, $\S$4.1].
The following lemmas give the height of a rational point and the relation between field extensions.
Lemma 3.5.
  1. (1)
    For $x\in X(K)$, $h_{\mathcal{L}_{\lambda}}(x)=\langle{\mathrm{deg}}(F_{P,x}),\lambda\rangle$.
  2. (2)
    If $x\in C_{w}(K)$, $F_{P,x}$ is of relative position $w$ with respect to $F_{Q}$.
  3. (3)
    If $F_{P}$ is in relative position $w$ with respect to $F_{Q}$, then $\langle{\mathrm{deg}}(F_{P}),\lambda\rangle\geq\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ for any antidominant character $\lambda$.
@@
  1. (1)
    By definition, $h_{\mathcal{L}_{\lambda}}(x)$ is the degree of $\sigma_{x}^{*}(\mathcal{L}_{\lambda})$ and $\langle{\mathrm{deg}}(F_{P,x}),\lambda\rangle$ is the degree of $F_{P,x}\times_{P}k_{\lambda}$. Then it follows from the equality
    $\sigma_{x}^{*}(\mathcal{L}_{\lambda})=\sigma_{x}^{*}(F\times_{P}k_{\lambda})=(\sigma_{x}^{*}F)\times_{P}k_{\lambda}=F_{P,x}\times_{P}k_{\lambda}$
  2. (2)
    We have a commutative diagram
    $\begin{array}[]{ccc}\sigma_{x}^{*}F&\longrightarrow&F=\coprod F_{Q}\times_{Q}QwP\\ \downarrow&&\downarrow\\ C&\stackrel{{\scriptstyle\sigma_{x}}}{{\longrightarrow}}&F/P=\coprod F_{Q}\times_{Q}QwP/P.\end{array}$
    So, the image lies in $F_{Q}\times_{Q}QwP$. The map $F_{Q}\times_{C}F_{P,x}\to C\times G$ factors as $F_{Q}\times_{C}F_{P,x}\to F_{Q}\times_{C}(F_{Q}\times_{Q}QwP)\to(F_{Q}\times_{C}F_{Q})\times_{Q}QwP$ and $(F_{Q}\times_{C}F_{Q})\times_{Q}QwP\to(C\times_{Q}Q)\times_{Q}QwP=C\times_{Q}QwP$.
  3. (3)
    See [10, Theorem4.1].
Proof of proposition 3.3.
For any finite extension $L/K$ and $x\in C_{w}(L)$, we define
  • $C_{L}$: normalization of $C$ in $L$
  • $X_{L}=X\times_{K}L$
  • $F_{C_{L}}=F\times_{C}C_{L}$
  • $\tilde{x}$: the corresponding morphism ${\rm Spec~}L\to X_{L}$, and define $\sigma_{\tilde{x}}$ similarly.
  • $\mathcal{L}_{C_{L},\lambda}$: the pullback of $\mathcal{L}_{\lambda}$.
By the above lemma(1),
$h_{\mathcal{L}_{\lambda}}(x)=\frac{1}{[L:K]}{\mathrm{deg}}\sigma_{\tilde{x}}^{*}\mathcal{L}_{\lambda}=\left\langle{\mathrm{deg}}(F_{C_{L},P,\tilde{x}}),\lambda\right\rangle$
Since $F$ admits strongly canonical reduction, the (strongly) canonical reduction of $F_{C_{L}}$ is $F_{C_{L},Q}=F_{Q}\times_{C}C_{L}$ and ${\mathrm{deg}}(F_{C_{L},Q})=[L:K]{\mathrm{deg}}(F_{Q})$ in $X(T)^{\vee}_{\mathbb{Q}}$. Then the proposition follows from the above lemma (2)(3). ∎
@@ 3.2. Height filtration and successive minima
Let $V$ be any $Q$-representation and let $V=\bigoplus_{\nu}V[\nu]$ be its weight decomposition. We define a filtration $V_{\bullet}$ on the vector space $V$: For any rational number $q\in\mathbb{Q}$, we define the subspace $V_{q}$ as the sum of weight spaces
@@
Then, by twisting the $Q$-subrepresentations $V_{q}$ above by $F_{Q}$, we obtain a filtration $V_{\bullet,F_{Q}}:=F_{Q}\times_{Q}V_{\bullet}$ of the vector bundle $V_{F_{Q}}=F_{Q}\times_{Q}V$ by subbundles
$0\neq V_{q_{1},F_{Q}}\subsetneq\cdots\subsetneq V_{q,F_{Q}}\subsetneq\cdots\subsetneq V_{q_{0},F_{Q}}=V_{F_{Q}}.$
Improving a result of Schieder [^1] The case of characteristic 0 and a specific case of positive characteristic were proven in [10], we proved that:
@@
It will suffice to check two assertions: all graded pieces are semistable vector bundles and their slopes are strictly decreasing.
As shown in [10, Lemma 5.1], first on $\mathrm{gr}_{q}V_{\bullet}$ the unipotent radical $U(L_{P})$ acts trivially, therefore $\mathrm{gr}_{q}V_{\bullet}$ is a $L_{Q}$-representation. Furthermore, we have
$\mathrm{gr}_{q}(V_{\bullet,F_{Q}})=V_{\bullet,F_{Q}}/V_{\bullet-1,F_{Q}}=(V_{\bullet}/V_{\bullet-1})_{F_{Q}}=(\mathrm{gr}_{q}V_{\bullet})_{F_{L_{Q}}}.$
If one writes the characters that appears in $V_{q}$ as $\sum{m_{i}}\lambda_{i}$, then its slope is computed as in
$\mu((\mathrm{gr}_{q}V_{\bullet})_{F_{L_{Q}}})=\frac{\sum\langle{\mathrm{deg}}F_{Q},\lambda_{i}\rangle m_{i}}{\sum m_{i}}=q.$
@@
To prove the semistability, consider the semisimplification $W$ of $\mathrm{gr}_{q}V_{\bullet}$ as a representation of $L_{Q}$}. From the same computation, we see that for each simple direct summand $W_{i}$ of $W$ the associated bundle $W_{i,F_{L_{Q}}}$ has slope $q$. Thus $(\mathrm{gr}_{q}V_{\bullet})_{F_{L_{Q}}}$ admits a filtration by subbundles such that each graded piece is a semistable vector bundle of slope $q$, where the semistability is obtained by applying Theorem 2.5 to ${L_{Q}}\to{\rm GL}(W_{i})$. Then $(\mathrm{gr}_{q}V_{\bullet})_{F_{M}}$ is also semistable of slope $q$ by the following. ∎
@@
If $E$ is a vector bundle on a curve $C$ that admits a filtration of subbundles
@@
and all graded pieces are semistable of slope equals to a given number $\lambda$, then $E$ is also semistable of slope equals to $\lambda$.
Proof.
The proof is elementary. See, for example [5, Page 580]
@@
The first assertion is obtained by taking $V$ to be $H^{0}\left(\overline{QwP}/P,M_{\lambda}\right)$ in Proposition 3.7. For the second assertion, it will suffice to show the highest weights in $H^{0}\left(\overline{QwP}/P,M_{\lambda}\right)$ belong to $w\lambda+\mathbb{Z}[\Delta_{Q}]$. The argument in [4] is also valid without any assumption on the characteristic of base field. Basically, one can use the restriction map
$\mathrm{H}^{0}(\overline{QwP}/P,M_{\lambda})\to\mathrm{H}^{0}(\overline{Bw^{\prime}P}/P,M_{\lambda})$
is surjective for all $w^{\prime}$ lying over the fibre of $w$ along $W/W_{P}\to W_{Q}\backslash W/W_{P}$, and the diagonal map
$\mathrm{H}^{0}(\overline{QwP}/P,M_{\lambda})\to\prod_{w^{\prime}}\mathrm{H}^{0}(\overline{Bw^{\prime}P}/P,M_{\lambda})$
is injective. These show that the highest weights appears and only appears in that of $\mathrm{H}^{0}(\overline{Bw^{\prime}P}/P,M_{\lambda})$ for all $w^{\prime}$, and they ahve the unique highest weights $w^{\prime}\lambda\in\lambda+\mathbb{Z}[\Delta_{Q}]$. ∎
@@
The arguments follows from the fact that the lower bound and essential minimum are invariant under taking closed subvariety as in [4, Proof of Theorem 2.1]. ∎
@@
$\tilde{X}=((\operatorname{Fr_{K}}^{n})^{*}F/P)_{K}$$X=(F/P)_{K}$${\rm Spec~}K$${\rm Spec~}K$$\scriptstyle{\phi_{K}}$$\scriptstyle{\operatorname{Fr}^{n}}$